so confused ..how would i simplify this expression:4R6S * 7R^3S^5 ??
Answer by sdmmadam@yahoo.com(530)

4R6S * 7R^3S^5
=(4*6*7)* (R*R^3)*(S*S^5) (grouping constants and quantities of the same base)
=168 * (R)^(1+3) * (s)^(1+5)
(when quantities of the same base are multiplied
write the base and add the powers.)
=168 (R)^4*(S)^6
Answer: 168[(R)^4*(S)^6]